OpenStudy (anonymous):

Solve the system of equations.4x + 7y = 60-4x + 7y = -4

OpenStudy (anonymous):

is that how it appears in the context

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it does not make sense 4x=7y cannot equal 60-4x=7y

OpenStudy (anonymous):

4x+7y=60 -4x+7y=-4

OpenStudy (anonymous):

okay that is good

OpenStudy (anonymous):

you have two equations and two unkowns you need to solve for x or y in one and them substitute that value in the other equation.

OpenStudy (anonymous):

i dont understand hopw to solve for one of them to do the other half

OpenStudy (anonymous):

4x+7y=60

OpenStudy (anonymous):

7y=60-4x

OpenStudy (anonymous):

y=60/7 - 4x/7

OpenStudy (anonymous):

i still dont understand at all.. this seriously makes no sense to me

OpenStudy (anonymous):

4x+7y =60 is balanced

OpenStudy (anonymous):

to keep the balance and solve for a variable you need to subtract 4x from both sides

OpenStudy (anonymous):

?????????????????

OpenStudy (anonymous):

(8,4)

OpenStudy (anonymous):

http://www.youtube.com/watch?v=iJQOTzzO124

OpenStudy (anonymous):

yes four and eight